Solve the quadratic equation. Show and explain each step in finding the solutions:
5x^2 + 2x = 3
This is what I have got when I have tried o do it
5x^2 + 2x = 3
Step 1: Write equation in standard form so one side equals 0 by adding -3 to both sides of equation
5x^2+2x-3=0
Step 2: Factor on left side of equation
(5x-1)(x+3) = 0
Step 3: Use zero-product rule
5x-1=0
x+3=0
Step 4: Solve equation x= 1/5 and x= -3


Found 2 solutions by stanbon, Edwin McCravy:
Answer by stanbon(75887)

5x^2 + 2x = 3
Step 1: Write equation in standard form so one side equals 0 by adding -3 to both sides of equation
5x^2+2x-3=0
======================
Factor:
5x^2 +5x-3x -3 = 0
--------------
5x(x+1)-3(x+1) = 0
(x+1)(5x-3) = 0
x = -1 or x = 3/5
=====================
Cheers,
Stan H.
===============

Answer by Edwin McCravy(20054)

Solve the quadratic equation. Show and explain each step in finding the solutions:
5x^2 + 2x = 3
This is what I have got when I have tried o do it
5x^2 + 2x = 3
Step 1: Write equation in standard form so one side equals 0 by adding -3 to both sides of equation
5x^2+2x-3=0
Step 2: Factor on left side of equation
(5x-1)(x+3) = 0

Here is your error. The correct factorization is

(5x-3)(x+1) = 0

Then the rest should have been:

Step 3: Use zero-product rule
5x-3=0
x+1=0
Step 4: Solve equation x= 3/5 and x= -1


Edwin